OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. P1 M P2 (3, 5) (-2, 0) ( , )

OpenStudy (amistre64):

distance moved from P1 to M, added to M

OpenStudy (amistre64):

M+(M-P) = 2M+P i wonder if thatd work

OpenStudy (amistre64):

2M-P that is

OpenStudy (amistre64):

2M= -4,0 -P =-3,-5 ---------- -7,-5 perhaps?

OpenStudy (anonymous):

should work

OpenStudy (amistre64):

should lol

OpenStudy (anonymous):

from 3 to -2 is left 5 units, left another 5 takes you to -7 from 5 to 0 is down 5 units, down another 5 take you to -5 (-7,-5)